Thursday afternoon                 Conference facilities rental                                         $425

Coffee-break services                                     $10 PER PERSON

Audiovisual equipment rental                                     $55

Evening reception                                           $25 PER PERSON

Friday                                      Conference facilities rental                                         $750

Coffee-break services, morning and               $15 PER PERSON

afternoon        

Continental Breakfast                                     $15/PERSON            

Luncheon                                                        $20 PER PERSON

Audiovisual equipment rental                                     $150

Saturday Afternoon                Conference facilities rental                                         $375

Continental Breakfast                                     $15 PER PERSON

                                                Coffee-break services                                     $10 PER PERSON

                                                Audiovisual equipment rental                                     $75

Program material and marketing would cost about $550. The association charged $130 for each participants last year and would like to use the same price this year.

a.Prepare a break-even chart for the conference and determine the break-even attendance level.

b. Suppose the association wanted to encourage student participation by charging a rate that would cover only the cost directly caused by their attendance. What price would you charge?

c. Prepare a budget for the event if you expect 110 people to attend (just show if there would be a surplus or deficit with 110 people attending).
Answer #1
Thursday afternoon                 Conference facilities rental 425
Coffee-break services 10 PER PERSON
Audiovisual equipment rental 55
Evening reception 25 PER PERSON
Friday                                      Conference facilities rental 750
Coffee-break services, morning and afternooon 15 PER PERSON
Continental Breakfast 15/PERSON            
Luncheon 20 PER PERSON
Audiovisual equipment rental 150
Saturday Afternoon                Conference facilities rental 375
Continental Breakfast 15 PER PERSON
                                                Coffee-break services 10 PER PERSON
                                                Audiovisual equipment rental 75

Total Rental Cost = 425+55+750+150+375+75 = $ 1830

Total fixed cost = Rental cost + Program material and marketing cost = 1830 + 550 = $ 2380

Variable cost per person = 10+25+15+15+20+15+10 = $ 110 per person

Profit = Attendance * (Fee per participant - Variable cost per person) - Fixed cost

a)

Break-even chart is as follows:

FORMULA: B2 =A2*(130-110)-2380 copy to B2:B20

Break-even attendance level = Fixed cost / (participant fee - variable cost) = 2380 / (130 - 110) = 119

b) That price would be equal to the variable cost per person = $ 110

c) With 110 people attending,

Total fee collection = 110*130 = $ 14,300

Total cost of the program = 2380 + 110*110 = $ 14,480

Total cost is greater than collection. Therefore, there will be a deficit.

Deficit = 14480 - 14300 = $ 180